OpenStudy (anonymous):

Evaluate 7a + 3b if a = -4 and b = -2.

OpenStudy (aravindg):

-34

OpenStudy (anonymous):

thank yu

OpenStudy (anonymous):

7(-4)+3(-2) = -28-6 =-34